# Finely Stratified Rerandomization Designs

### v1

@@ 6 Linear Adjustment
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.2. Then we have $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V(\alpha))$ with $V(\alpha)=v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\alpha^{\prime}w|\psi)]$ and $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow\mathcal{N}(0,V_{\phi}+V(\alpha))$.
@@
$\alpha_{0}\in\operatorname*{argmin}_{\alpha\in\mathbb{R}^{d_{w}\times d_{g}}}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\alpha^{\prime}w|\psi)].$ (6.1)
Note that if $w=h$ then $\alpha_{0}=\Gamma_{0}$, as in Theorem 3.4. If $E[\operatorname{Var}(w|\psi)]\succ 0$, then the unique minimizer of Equation 6.1 is the partially linear regression coefficient $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi\delta(W,\theta_{0})|\psi)]$. The main result of this section shows that adjustment by a consistent estimate of $\alpha_{0}$ restores asymptotic normality.
Theorem 6.3 (Adjusted GMM).
@@
$V_{\phi}=\operatorname{Var}(\Pi\phi(W,\theta_{0}))\quad\quad V_{\delta}^{a}=\min_{\alpha\in\mathbb{R}^{d_{w}\times d_{g}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\alpha^{\prime}w|\psi)].$
@@
Remark 6.5 (One Step Adjustment).
In many cases, the optimal coefficient $\alpha_{0}$ may not depend on $\theta_{0}$ at all, allowing optimal adjustment to be done in one step. For instance, whenever $\delta(W,\theta)=u(\psi,\theta)+v(W)$ for some functions $u,v$, then $\alpha_{0}$ does not depend on $\theta_{0}$. This happens in Example 3.9, where we have $\delta(W,\theta)=(1-p)Y(1)+pY(0)=\bar{Y}$. The optimal adjustment coefficient is given by $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\bar{Y}|\psi)]$.
Theorem 6.3 that ex-post linear adjustment can be used to remove the non-Gaussian component of the asymptotic distribution that arises due to rerandomization. In the next section, we exploit this result to provide standard t-statistic based confidence intervals for $\theta_{n}$ and $\theta_{0}$.
@@ 7.1 Asymptotically Exact Inference
To define our variance estimator, we begin with some definitions. Let $\mathcal{G}_{n}$ denote the set of groups constructed in Definition 2.1. For each $g\in\mathcal{G}_{n}$ define the centroid $\bar{\psi}_{g}=|g|^{-1}\sum_{i\in g}\psi_{i}$. Let $\nu:\mathcal{G}_{n}\to\mathcal{G}_{n}$ be a bijective matching between groups satisfying $\nu(g)\not=g$, $\nu^{2}=\operatorname{Id}$, and the homogeneity condition
$\frac{1}{n}\sum_{g\in\mathcal{G}_{n}}|\bar{\psi}_{g}-\bar{\psi}_{\nu(g)}|_{2}^{2}=o_{p}(1).$ (7.1)
@@
$\widehat{v}_{1}$ =$n^{-1}\sum_{g\in\mathcal{G}^{\nu}_{n}}\frac{1}{a(g)-1}\sum_{i\not=j\in g}\widehat{m}_{i}\widehat{m}_{j}^{\prime}D_{i}D_{j}/p$
$\widehat{v}_{0}$ =$n^{-1}\sum_{g\in\mathcal{G}^{\nu}_{n}}\frac{1}{(k-a)(g)-1}\sum_{i\not=j\in g}\widehat{m}_{i}\widehat{m}_{j}^{\prime}(1-D_{i})(1-D_{j})/(1-p)$
$\widehat{v}_{10}$ $=n^{-1}\sum_{g\in\mathcal{G}_{n}}\frac{k}{a(k-a)}(g)\sum_{i,j\in g}\widehat{m}_{i}\widehat{m}_{j}^{\prime}D_{i}(1-D_{j}).$
@@
By Theorem 6.3, $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow N(0,V_{\phi}+V_{\delta}^{a})$. Then the variance estimation result above allows for joint inference on $\theta_{0}$ using e.g. standard Wald-test or t-statistic based confidence regions.
@@ 7.2 Inference on the Finite Population Parameter
As noted above, the asymptotic variance $V_{\delta}^{a}$ in Theorem 6.3 for estimating the finite population parameter $\theta_{n}$ is generically not identified. This happens because it depends on terms of the form $\operatorname{Var}(\delta|\psi)\propto\operatorname{Var}(g_{1}|\psi)+\operatorname{Var}(g_{0}|\psi)-2\operatorname{Cov}(g_{1},g_{0}|\psi)$, with $g_{d}=g(d,X,S(d),\theta_{0})$. However, $S(1)$ and $S(0)$ are never simultaneously observed (Neyman, (1990)), so $\operatorname{Cov}(g_{1},g_{0}|\psi)$ is generically not identified. We work with linear contrasts $c^{\prime}\theta_{n}$ since this allows us to tighten our upper bounds on the (non-identified) variance. To do so, let $\widehat{u}_{1}=E_{n}[\frac{D_{i}}{p}\widehat{m}_{i}\widehat{m}_{i}^{\prime}]-\widehat{v}_{1}$ and $\widehat{u}_{0}=E_{n}[\frac{1-D_{i}}{1-p}\widehat{m}_{i}\widehat{m}_{i}^{\prime}]-\widehat{v}_{0}$ using the estimator components above and consider the variance estimator
$\widehat{V}_{\delta}(c)=v_{D}([c^{\prime}\widehat{u}_{1}c]^{1/2}+[c^{\prime}\widehat{u}_{0}c]^{1/2})^{2}.$ (7.3)
By Theorem 6.3, we have $\sqrt{n}(c^{\prime}\widehat{\theta}_{adj}-c^{\prime}\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,c^{\prime}V_{\delta}^{a}c)$. Our next result shows how to consistently estimate an upper bound on this asymptotic variance.
@@
The variance upper bound $\bar{V}_{\delta}(c)\geq c^{\prime}(V_{\phi}+V_{\delta}^{a})c$, so the confidence intervals derived from this approach are always weakly shorter than those using the variance estimator in Equation 7.2. See Section 8.6 in the appendix for an explicit comparison. The upper bound $\bar{V}_{\delta}(c)$ incorporates the efficiency gains from stratification, rerandomization, and adjustment. However, this upper bound is generally not sharp (Aronow et al., (2014)). We leave sharp upper bounds on the asymptotic variance matrix $V_{\delta}^{a}$ to future work.

### v2

@@ 6 Restoring Normality
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.2. Then we have $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V(\alpha))$ with $V(\alpha)=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\alpha^{\prime}w|\psi)]$ and $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow\mathcal{N}(0,V_{\phi}+V(\alpha))$.
@@
$\alpha_{0}\in\operatorname*{argmin}_{\alpha\in\mathbb{R}^{d_{w}\times d_{g}}}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\alpha^{\prime}w|\psi)].$ (6.1)
Note that if $w=h$ then $\alpha_{0}=\gamma_{0}$, as in Theorem 3.5. If $E[\operatorname{Var}(w|\psi)]\succ 0$, then the unique minimizer of Equation 6.1 is the partially linear regression coefficient $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi a(W,\theta_{0})|\psi)]$. Observe that the optimal adjustment coefficient $\alpha_{0}$ varies with the stratification variables $\psi$, as observed in Cytrynbaum, (2024) and Bai et al., (2023) for ATE estimation. The main result of this section shows that adjustment by a consistent estimate of $\alpha_{0}$ restores asymptotic normality.
Theorem 6.3 (Restoring Normality).
@@
$V_{\phi}=\operatorname{Var}(\Pi\phi(W,\theta_{0}))\quad\quad V_{a}^{adj}=\min_{\alpha\in\mathbb{R}^{d_{w}\times d_{g}}}v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\alpha^{\prime}w|\psi)].$
@@
In some cases, $\alpha_{0}$ is independent of $\theta_{0}$. For example, if $a(W,\theta)=a_{1}(\psi,\theta)+a_{2}(W)$ then $\alpha_{0}=E[\operatorname{Var}(w|\psi)]^{-1}E[\operatorname{Cov}(w,\Pi a_{2}(W)|\psi)]$ does not depend on $\theta_{0}$. In such cases, one-step optimal adjustment is possible.
Corollary 6.5 (One-step Adjustment).
Suppose $a(W,\theta)=a_{1}(\psi,\theta)+a_{2}(W)$. Then for any $\theta\in\Theta$, substituting $g_{i}=g(D_{i},R_{i},S_{i},\theta)$ for $\widehat{g}_{i}$ in $\widehat{\alpha}$ above, we have $\widehat{\alpha}=\alpha_{0}+o_{p}(1)$.
One-step adjustment is possible in many linear GMM problems, including the best linear predictor of treatment effects parameter in Example 3.11.
Example 6.6 (Treatment Effect Heterogeneity).
Continuing Example 3.11 with score $g(Y,D,X,\theta)=(HY-X^{\prime}\theta)X$ and $\theta_{n}=\operatorname*{argmin}_{\theta}E_{n}[(\tau_{i}-X_{i}^{\prime}\theta)^{2}]$, recall that $a(W,\theta_{0})=\bar{Y}X$ and $\Pi=E[XX^{\prime}]^{-1}$. Letting $\theta=0$ gives $g(Y,D,X,0)=HYX$. By Corollary 6.5, we have $\widehat{\alpha}=\alpha_{0}+o_{p}(1)$ for the adjustment coefficient matrix
$\widehat{\alpha}=v_{D}E_{n}[\check{w}_{i}\check{w}_{i}^{\prime}]^{-1}E_{n}[H_{i}^{2}Y_{i}\check{w}_{i}X_{i}^{\prime}]E_{n}[X_{i}X_{i}^{\prime}]^{-1}.$ (6.2)
This allows us to estimate treatment effect heterogeneity relative to a low-dimensional vector of important covariates $X$, adjusting optimally for a larger set of covariates $w$ ex-post in order to improve precision, as well as restore asymptotic normality when $A\not=\mathbb{R}^{d_{h}}$. In the case $X=1$ (SATE estimation), $\widehat{\alpha}$ is equivalent to the “tyranny-of-the-minority” style estimator proposed in Cytrynbaum, (2023).
@@ 7.1 Asymptotically Exact Inference
To define our variance estimator, we begin with some definitions. Let $\mathcal{S}_{n}$ denote the set of groups constructed in Definition 2.1. For each $s\in\mathcal{S}_{n}$ define the centroid $\bar{\psi}_{s}=|s|^{-1}\sum_{i\in s}\psi_{i}$. Let $\nu:\mathcal{S}_{n}\to\mathcal{S}_{n}$ be a bijective matching between groups satisfying $\nu(s)\not=s$, $\nu^{2}=\operatorname{Id}$, and the homogeneity condition
$\frac{1}{n}\sum_{s\in\mathcal{S}_{n}}|\bar{\psi}_{s}-\bar{\psi}_{\nu(s)}|_{2}^{2}=o_{p}(1).$ (7.1)
@@
$\widehat{v}_{1}$ =$n^{-1}\sum_{s\in\mathcal{S}^{\nu}_{n}}\frac{1}{a(s)-1}\sum_{i\not=j\in s}\widehat{m}_{i}\widehat{m}_{j}^{\prime}D_{i}D_{j}/p$
$\widehat{v}_{0}$ =$n^{-1}\sum_{s\in\mathcal{S}^{\nu}_{n}}\frac{1}{(k-a)(s)-1}\sum_{i\not=j\in s}\widehat{m}_{i}\widehat{m}_{j}^{\prime}(1-D_{i})(1-D_{j})/(1-p)$
$\widehat{v}_{10}$ $=n^{-1}\sum_{s\in\mathcal{S}_{n}}\frac{k}{a(k-a)}(s)\sum_{i,j\in s}\widehat{m}_{i}\widehat{m}_{j}^{\prime}D_{i}(1-D_{j}).$
@@
By Theorem 6.3, $\sqrt{n}(\widehat{\theta}_{adj}-\theta_{0})\Rightarrow N(0,V_{\phi}+V_{a}^{adj})$. Then the variance estimation result above allows for joint inference on $\theta_{0}$ using e.g. standard Wald-test or t-statistic based confidence regions.
@@ 7.2 Inference on the Finite Population Parameter
As noted above, the asymptotic variance $V_{a}^{adj}$ in Theorem 6.3 for estimating the finite population parameter $\theta_{n}$ is generically not identified. This happens because it depends on terms of the form $\operatorname{Var}(a|\psi)\propto\operatorname{Var}(g_{1}|\psi)+\operatorname{Var}(g_{0}|\psi)-2\operatorname{Cov}(g_{1},g_{0}|\psi)$, with $g_{d}=g(d,X,S(d),\theta_{0})$. However, $S(1)$ and $S(0)$ are never simultaneously observed (Neyman, (1990)), so $\operatorname{Cov}(g_{1},g_{0}|\psi)$ is generically not identified. We work with linear contrasts $c^{\prime}\theta_{n}$ since this allows us to tighten our upper bounds on the (non-identified) variance. To do so, let $\widehat{u}_{1}=E_{n}[\frac{D_{i}}{p}\widehat{m}_{i}\widehat{m}_{i}^{\prime}]-\widehat{v}_{1}$ and $\widehat{u}_{0}=E_{n}[\frac{1-D_{i}}{1-p}\widehat{m}_{i}\widehat{m}_{i}^{\prime}]-\widehat{v}_{0}$ using the estimator components above and consider the variance estimator
$\widehat{V}_{a}(c)=v_{D}([c^{\prime}\widehat{u}_{1}c]^{1/2}+[c^{\prime}\widehat{u}_{0}c]^{1/2})^{2}.$ (7.3)
By Theorem 6.3, we have $\sqrt{n}(c^{\prime}\widehat{\theta}_{adj}-c^{\prime}\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,c^{\prime}V_{a}^{adj}c)$. Our next result shows how to consistently estimate an upper bound on this asymptotic variance.
@@
The variance upper bound $\bar{V}_{a}(c)\geq c^{\prime}(V_{\phi}+V_{a}^{adj})c$, so the confidence intervals derived from this approach are always weakly shorter than those using the variance estimator in Equation 7.2. See Section 8.8 in the appendix for an explicit comparison. The upper bound $\bar{V}_{a}(c)$ incorporates the efficiency gains from stratification, rerandomization, and adjustment. However, this upper bound is generally not sharp (Aronow et al., (2014)). We leave sharp upper bounds on the asymptotic variance matrix $V_{a}^{adj}$ to future work.